March 24, 1991

Aspects of Quantum Field Theory in Curved Space-Time

S. A. Fulling

Corrections for second printing

Chapter Page Line  Old  New

2 21 10–11 equation of motion of other fields, which may equations of motion of other fields; those equations may

2  23 20  = D_{j}b^{j}  = \frac{1}{2}D_{j}b^{j}
2  25 15  (nπ)^{2},  (nπ)^{2} (n a positive integer),

2 26

−2–−1

(pointwise), H is semibounded.

(as a function), H is semibounded (as an operator).

2  29 14  λ^{2}(σ  L^{2}(σ
2  29 15  λ^{2}(σ  L^{2}(σ
2  40 7  (H − z)^{−1}u  (H − z)u
2  47 −9  w_{k}  A_{k}
3  61 14  to e_{j_{1}} ⊗ ⋯ ⊗ e_{j_{n}}  to e_{j_{1}} ⊙ ⋯ ⊙ e_{j_{n}}
3  66 (3.21)  e^{+iω_{j}}  e^{+itω_{j}}
3  68 13  ∇ϕ(t, x)  ∇ϕ(t, x), ∂_{t}ϕ(t, x)

3 68 18 explicitly in the integral defining L.) in the elements of integration, but in L itself.)

3  68 19  ϕ̇  ϕ̇*
4  75 5  e^{−iωs}  e^{iωs}
4  75 7  e^{−iωs}  e^{iωs}
4  80 −11  AB − BA  AB + BA
4  80 −5  G  iG

4  80 −3  G  iG
4  84 14  w_{j}^{2p+1}  2ω_{j}^{2p+1}
4  84 15  w_{j}^{2p+1}  2ω_{j}^{2p+1}
4  86 −4  −e  −ϵ
5  96 1  have f  have ϕ
5  101 −5  iz  iζ
6  118 −3  L\sqrt{g}  L
6  130 4  \underline{γ}  −\underline{γ}
7  135 −11  \frac{1}{2}R  \frac{1}{2}Rg_{μν}
7  138 (7.5)  |α_{k}| − |β_{k}|  |α_{k}|^{2} − |β_{k}|^{2}
8  172 (8.19)  = R  = −R

8 173 11 distinct ordered pairs (μν) choices of pairs (μν) of canonically ordered distinct indices

8  175 −8  ẋ^{ν}ẋ^{ν}  ẋ^{μ}ẋ^{ν}
9  208 4  which  that
9  208 5  quartic  inversely quartic
9  208 6  r  ρ
9  208 7  quadratic  inversely quadratic
9  208 8  term linear  term inversely linear